There are 5 computers in a store. Among them, 3 are brand new, and 2 are
refurbished. How many ways are there to choose 3 computers that include 1
refurbished computer!

Found 2 solutions by saw, Edwin McCravy:
Answer by saw(34)

Let b = brand new
Let r = refurbished
3b+2r = 5
b+b+b+r+r = 5
b+b+r
b+r+b
r+b+b
nPr = n!/(n-r)! = 3!/(3-1)! = 3!/2! = 6/2 = 3
There are 3 ways to choose 3 computers that include 1 refurbished computer.






Answer by Edwin McCravy(20054)

There are 5 computers in a store. Among them, 3 are brand new, and 2 are
refurbished. How many ways are there to choose 3 computers that include 1
refurbished computer!
Choose the 1 unfurbished one from the 2.

(2 refurbished CHOOSE 1) = C(2,1) = (2∙1)/1! = 2 ways.

For each of those wats, we choose the 2 brand new ones from the 3.

(3 brand new ones CHOOSE 2) = C(3,2) = (3∙2∙1)/2! = 6/2 = 3 ways.

We multiply (because it's an "AND" situation).

2∙3 = 6 ways.

Checking:

If the 3 brand new ones are A,B, and C and the 2 refurbished ones are X and Y,
we can choose:

1. A,B,X
2, A,B,Y
3. A,C,X
4. A,C,Y
5. B,C,X
6. B.C.Y

Edwin

Edwin